100 grams of sodium chloride react with 100 grams of silver nitrate. Limiting Reactant? NaCl+AgNO 3 NaNO 3 +AgCl


1
Expert's answer
2021-03-26T05:45:40-0400

Molar Mass of NaCl = 58.44 g/mol

Molar Mass of AgNO3 = 169.87 g/mol

100g/58.44g = 1.711 M of NaCl

100/169.87 = 0.589 M of AgNO3

ratio = 1:1

= 1.711 ×1 mol AgNO3 = 1.711 mol AgNO3

Because we have only 0.589M of AgNO3 compared to 1.711M needed, the we have less AgNO3 for reacting NaCl. Therefore, AgNO3 is the limiting reactant.